# From Consistency to Complementarity: Aligned and Disentangled Multi-modal Learning for Time Series Understanding and Reasoning

### v1

@@ 5.1 Experimental Setups
Baseline. We compare our model against three types of baselines, grouped by the formulation of time series modalities: (1) Numerical-centric models directly process time series values, including general LLMs (GPT-4o (Hurst et al., 2024), Qwen3 (Yang et al., 2025), DeepSeek-V3.2 (Liu et al., 2025a), Gemini 3 Pro (Google, 2025), GPT-5.2 (OpenAI, 2025)) and time series-specific MLLMs (ChatTime (Wang et al., 2025a), ChatTS (Xie et al., 2025), ITFormer (Wang et al., 2025b), InstructTime (Cheng et al., 2025)); (2) Visual-centric models leverage time series plots (GPT-4o, Qwen3-VL (Shuai et al., 2025), Gemini 3 Pro, GPT-5.2); and (3) Numerical+Visual models integrate both modalities (GPT-4o, Qwen3-VL, Gemini 3 Pro, GPT-5.2, GEM (Lan et al., 2025)). General-purpose LLMs/MLLMs are accessed via APIs, while time-series-specialized MLLMs are fine-tuned on unified LLM backbones and training data. More details are provided in Appendix B.2.

### v2

@@ 5.1 Experimental Setups
Baseline. We compare our model against three types of baselines, grouped by the formulation of time series modalities: (1) Numerical-centric models directly process time series values, including general LLMs (GPT-4o (Hurst et al., 2024), Qwen3 (Yang et al., 2025), DeepSeek-V3.2 (Liu et al., 2025a), Gemini 3 Pro (Google, 2025), GPT-5.2 (OpenAI, 2025)) which formulate numerical values into textual prompts, and time series-specific MLLMs (ChatTime (Wang et al., 2025a), ChatTS (Xie et al., 2025), ITFormer (Wang et al., 2025b), InstructTime (Cheng et al., 2025)) which introduce specialized time series encoders; (2) Visual-centric models leverage time series plots (GPT-4o, Qwen3-VL (Shuai et al., 2025), Gemini 3 Pro, GPT-5.2); and (3) Numerical+Visual models integrate both modalities (GPT-4o, Qwen3-VL, Gemini 3 Pro, GPT-5.2, GEM (Lan et al., 2025)). General-purpose LLMs/MLLMs are accessed via APIs, while time-series-specialized MLLMs are fine-tuned on unified LLM backbones and training data. More details are provided in Appendix B.2.